A rotating wheel requires 5.00 s to rotate 26.0 revolutions. Its angular velocity at the end of the 5.00 s interval is 95.0 rad/s. What is the constant angular acceleration of the wheel?
_________ rad/s2
Answer #1

Since angular acceleration is constant, therefore we use equations of motion for angular motion,

1) w​​​​​f = w​​​​​​I + \alpha t

wf - final angular velocity

wi - initial angular velocity

\alpha - angular acceleration

t - time taken

2) \theta = w​​​​​​i​​​t + 1/2 \alpha t2

\theta - angular displacement

t= 5s

w​​​​​​F = 95 rad/s

\theta= no. Of revolution x 2π = 26x2π = 52π

Putting in equations,

1) 95 = wi + \alpha (5)

=> Wi = 95-5a

2) 52π = 5Wi + (1/2)(25)\alpha

Using 1

52π = 5(95-5a) + 12.5a

=> 12.5a = 475-52(3.14) = 311.72

=> \alpha = 311.72/12.5 = 24.93 rad/s​​​​​​2